OpenStudy (anonymous):

solve the equation by graphing : y=6 y=4/3x+2

OpenStudy (anonymous):

no idea how to do this.

hero (hero):

In order to solve this problem, you'll need to know the following: 1. How to plot lines on an xy plane 2. How find the intersection point of two lines 3. How to plot points on an xy plane Which of the following are you having trouble with?

OpenStudy (anonymous):

In this instance, you would graph the line y = 6 as well as y = 4/3x+2. The point of intersection should be the solution.

OpenStudy (anonymous):

@hero 2 and 3

hero (hero):

So, what you're telling me is you've plotted both lines on the xy plane, however, you haven't noticed that they intersect, correct?

OpenStudy (anonymous):

how do i plot the second one?

OpenStudy (anonymous):

In graphing, the intersection is when the two lines cross. If you correctly graph the lines and the intersection point is an integer, you can find where they intersect. to graph the second equation, when x=0, y = 2. So the y intercept is (0,2). That's one point. To plot the rest of the line, you go over 3, up 4, plot point, connect etc.

hero (hero):

wombat, I know you're trying to help, but it is less confusing if one person tries to help her at a time.

OpenStudy (anonymous):

@hero so for the first equation i plot (0,6)? correct

hero (hero):

Just to clarify, y = 6 is a line, not a point. When you plot y = 6, you consider all the points on an xy plane that have the coordinate (x,6)

OpenStudy (anonymous):

so its a horizontal line crossing the axis at (0,6)?

hero (hero):

Yes, correct

OpenStudy (anonymous):

okay so about the other equation how do i do that one??

hero (hero):

To plot the other line, you'll need to plot two points that represent the x-intercept and the y-intercept of that line. In other words, one of the points will intersect the x-axis. The other point will intersect the y-axis. To find the x-intercept, set y = 0 then solve for x To find the y-intercept, set x= 0, then solve for y

OpenStudy (anonymous):

so 0=4/3x+2 is x intercept ?

hero (hero):

For the x-intercept, you'll end up with some point (x,0) For the y-intercept, you'll end up with some point (0,y)

hero (hero):

jenny, when I say "solve for x" I mean to isolate x on one side of the equation such that it is by itself on one side of the equation. We need an equation where x = some number You'll need to use your knowledge of algebra to manipulate that equation to get x alone.

hero (hero):

If you need help doing this, please say something.

OpenStudy (anonymous):

yes i need help

hero (hero):

Jenni, you've never solved an equation for x before?

hero (hero):

Suppose you had 3x + 5 = 17 and you had to find x, what would you do?

OpenStudy (anonymous):

Do you still need help?

hero (hero):

I'm still helping her.

OpenStudy (anonymous):

Ok Ok Ok.

OpenStudy (anonymous):

you take 5 from both sides to get 3x=12 then divide both sides to get x by itself to get x=4